# Graphon Quantum Filtering systems

### v1

@@ 5.2 Propagation of chaos
[…]
To address the mean-field limit, we extend the method introduced by Pickl [60] and later adapted to the stochastic setting in [41, 43], to the case of multi-class bosons system.
For each class, we consider $N$ independent copies, each driven by its own Wiener process. Denote by $\gamma_{j_{l^{\prime}}}$ the $l^{\prime}$-th particle in class $j$, whose dynamics follow the SDE:
$\mathrm{d}\gamma_{t,j_{l^{\prime}}}$ $=-\mathrm{i}\Big{[}\tilde{H}+\frac{1}{c}\sum_{i=1}^{c}w_{ji}A^{\mathbb{E}[\gamma_{t,i}]},\gamma_{t,j_{l^{\prime}}}\Big{]}\mathrm{d}t+\big{(}L\gamma_{t,j_{l^{\prime}}}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,j_{l^{\prime}}}\}\big{)}\mathrm{d}t$
@@
The goal is to introduce a quantity that detects if $\boldsymbol{\rho}^{cN}_{t}$ is tensorized in “good way” each times. To do this, it is sufficient to show that the number of ”bad” particles, that is particles deviating from the desired state $\boldsymbol{\gamma}_{i_{l}}$, is small.
In order to ascertain the ratio of bad particles, we take advantage of counting operator running through all possible combinations of bad particles of each class. Let’s introduce an operator which counts the number of bad particles in each class:
@@
$\mathbf{P}_{\hat{n}}$ $:=\prod_{i=1}^{c}\mathbf{P}_{i,n_{i}},\;\hat{n}=\begin{pmatrix}n_{1}\\ \vdots\\ n_{c}\end{pmatrix}.$
Consequently we can now define a function that quantifies the number of bad particles in the system by counting all weighted combination of numbers of bad particles.
For any sub-additive function $\varpi:\{0,\dots,cN\}\to\mathbb{R}_{+}$, we define
@@
that can be equivalently written as
@@
Moreover, if we count in an unbiased way, i.e., $\varpi(k)=\frac{k}{cN}$, then:
$\mathbb{E}\Big{[}\mathcal{E}_{N,\varpi}(t)\Big{]}\leq\frac{1}{c}\sum_{i=1}^{c}\mathbb{E}\Big{[}\mathcal{D}^{i_{l}}_{N}(t)\Big{]},\,\;\forall l\in[N],$
@@
=$\stackrel{{\scriptstyle\text{Fubini}}}{{=}}\langle\boldsymbol{\Psi}_{t}^{cN},\sum_{i=1}^{c}\sum_{n_{1},\dots,n_{c}=0}^{N}\Big{(}\varpi(n_{i})\mathbf{P}_{i,n_{i}}\prod_{j\neq i}^{c}\mathbf{P}_{j,n_{j}}\Big{)}\boldsymbol{\Psi}_{t}^{cN}\rangle$
@@
$\mathbb{E}\Big{[}\mathcal{E}_{N,\varpi}^{i}(t)\Big{]}$ $=\sum_{n_{i}=0}^{N}\mathbb{E}\Big{[}\langle\boldsymbol{\Psi}_{t}^{cN},\frac{n_{i}}{cN}\mathbf{P}_{i,n_{i}}\boldsymbol{\Psi}_{t}^{cN}\rangle\Big{]},$
$=\frac{1}{c}\mathbb{E}\Big{[}\langle\boldsymbol{\Psi}_{t}^{cN},\frac{1}{N}\sum_{n_{i}=0}^{N}{n_{i}}\mathbf{P}_{i,n_{i}}\boldsymbol{\Psi}_{t}^{cN}\rangle\Big{]},$
$=\frac{1}{c}\mathbb{E}\Big{[}\langle\boldsymbol{\Psi}_{t}^{cN},\frac{1}{N}\sum_{n_{i}=0}^{N}\sum_{l=1}^{N}{\boldsymbol{g}_{t,i_{l}}}\mathbf{P}_{i,n_{i}}\boldsymbol{\Psi}_{t}^{cN}\rangle\Big{]},$
=$\stackrel{{\scriptstyle\text{Invariance in law }}}{{=}}\frac{1}{c}\mathbb{E}\Big{[}\langle\boldsymbol{\Psi}_{t}^{cN},\boldsymbol{g}_{t,i_{l}}\boldsymbol{\Psi}_{t}^{cN}\rangle\Big{]},\;\;\forall l\in[N],$
$=\frac{1}{c}\mathbb{E}\Big{[}\mathcal{D}_{N}^{i_{l}}(t)\Big{]},$
@@
The above lemma is crucial because it allows us, to compute the numbers of bad particles class by class, and to work directly with the trace norm. Indeed, the quantity $\mathcal{D}_{N}^{i_{l}}(t)$ can be seen as a trace-norm distance between the state of particle $i_{l}$ in the $cN$-particles system and the corresponding limiting state in the limiting graphon system, i.e., $\mathcal{D}^{i_{l}}_{N}(t)=\big{\|}\rho_{t,i_{l}}-\gamma_{t,i_{l}}\big{\|}_{1}$.
We now state some technical lemmas that are useful to establish our result. The following lemma results from a slight modification of [39, Lemma 3.10] and [43, Lemma A.1].
Lemma 3.
Let $f:\mathbb{N}\to\mathbb{R}$ be a function and define the operator $\hat{f}_{i}=\sum_{k}f(k)\mathbf{P}_{i,k}$. Consider two products $\boldsymbol{Q}_{r_{i},r_{j}}^{(a)}$, $\boldsymbol{Q}_{r_{i},r_{j}}^{(b)}$ defined in the following way :
@@
$\|a\|^{2}_{2}$ $=\int_{\mathfrak{X}^{4}}|a(x,y;x^{\prime},y^{\prime})|^{2}\mu(\mathrm{d}x)\mu(\mathrm{d}y)\mu(\mathrm{d}x^{\prime})\mu(\mathrm{d}y^{\prime})<\infty$
@@
$(\gamma_{2}$ $A_{12}\gamma_{2})f(x_{1},x_{2})=\int_{\mathfrak{X}}\gamma_{2}(x_{2},y_{2})(A_{12}\gamma_{2}f)(x_{1},y_{2})\mu(\mathrm{d}y)$
@@
$=\int_{\mathfrak{X}^{2}}\Big{(}\int_{\mathfrak{X}^{2}}\gamma(x_{2},y_{2})a(x_{1},y_{2};x^{\prime}_{1},z_{2}){\gamma(z_{2},x^{\prime}_{2})}\mu(\mathrm{d}y_{2})\mu(\mathrm{d}z_{2})\Big{)}f(x_{1}^{\prime},x_{2}^{\prime})\mu(\mathrm{d}x_{1}^{\prime})\mu(\mathrm{d}x_{2}^{\prime})$
@@
We are now ready to state the second main result of the paper.
Theorem 3 (Propagation of chaos estimate).
@@
$\mathbb{E}[\mathcal{E}_{N,\hat{f}}(t)]$ $\leq e^{Ct}\Big{(}\mathcal{E}_{N,\hat{f}}(0)+\|{w}^{G_{cN}}-{w}\|_{\square}+\frac{1}{\sqrt{N}}\Big{)},$
@@
We proceed ”à la Grönwall”. Recall that in unbiased counting, we have
$\mathbb{E}[\mathcal{E}_{N,\hat{f}}(t)]\leq\frac{1}{c}\sum_{j=1}^{c}\mathbb{E}[\mathcal{D}^{j_{l}}_{N}(t)],\quad\forall l\in[N].$
@@
where $\mathcal{P}_{1}^{j}$ collects the drift terms that are independent of $L_{j_{1}}$ and $\mathcal{P}_{2}^{j}$ collects the drift terms that depend on $L_{j_{1}}$.
The estimates used to bound the quantity $|\mathcal{P}_{2}^{j}|$ are found in the lemmas of [41, Appendix A], and the extensions for all density operators provided in [26]. We have
$\mathcal{P}_{2}^{j}$ $=\mathrm{t}\mathrm{r}\Big{(}\Big{(}\frac{1}{2}L_{j}^{\dagger}L_{j}\boldsymbol{\rho}_{t}^{cN}+\frac{1}{2}\boldsymbol{\rho}_{t}^{cN}L_{j}^{\dagger}L_{j}-L_{j}\boldsymbol{\rho}_{t}^{cN}L_{j}^{\dagger}\Big{)}\boldsymbol{\gamma_{t,j}}+\boldsymbol{\rho}_{t}^{cN}\Big{(}\frac{1}{2}L_{j}^{\dagger}L_{j}\boldsymbol{\gamma}_{t,j}+\frac{1}{2}\boldsymbol{\gamma}_{t,j}L^{\dagger}_{j}L_{j}-L_{j}\boldsymbol{\gamma}_{t,j}L_{j}^{\dagger}\Big{)}$
@@
$\mathcal{P}_{1}^{j}$ $=\mathrm{i}\,\mathrm{t}\mathrm{r}\Big{(}\boldsymbol{\rho}_{t}^{cN}[\tilde{\mathbf{H}}_{j},\boldsymbol{\gamma}_{t,j}]+\boldsymbol{\rho}_{t}^{cN}[\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}[\boldsymbol{\gamma_{t,i}}]},\boldsymbol{\gamma}_{t,j}]\Big{)}$
@@
$\mathcal{P}_{1}^{j}$ $=\mathrm{i}\,\mathrm{t}\mathrm{r}\Big{(}\big{[}\tilde{\mathbf{H}}_{j},\boldsymbol{\gamma}_{t,j}\big{]}\boldsymbol{\rho}_{t}^{cN}+[\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}[{\gamma_{t,i}}]},\boldsymbol{\gamma}_{t,j}]\boldsymbol{\rho}_{t}^{cN}\Big{)}$
@@
All terms with $\tilde{\mathbf{H}}_{i_{l}}$ vanish because $[\tilde{\mathbf{H}}_{i_{l}},\boldsymbol{g}_{t,j}]=0$ when $i_{l}\neq j_{1}$ [^1] This follows since the operators $\mathbf{\tilde{H}}_{i_{l}},\boldsymbol{g}_{t,j}$ act on different variables. and the two remaining $\tilde{\mathbf{H}}_{j}$ terms cancel each other,
$\mathcal{P}_{1}^{j}$ $=\mathrm{i}\,\mathrm{t}\mathrm{r}\Big{(}\Big{[}\frac{1}{2cN}\sum_{i_{l}=1_{1}}^{c_{N}}\sum_{j_{l^{\prime}}=i_{l}}^{c_{N}}\xi_{j_{l^{\prime}}i_{l}}A_{j_{l^{\prime}}i_{l}},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}-\mathrm{i}\,\mathrm{t}\mathrm{r}\Big{(}\Big{[}\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}[\boldsymbol{\gamma_{t,i}}]},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}$
$=\mathrm{i}\,\mathrm{t}\mathrm{r}\Big{(}\Big{[}\frac{1}{2cN}\sum_{i_{l}=1_{1}}^{c_{N}}\sum_{h_{k}=i_{l}}^{c_{N}}\xi_{j_{l^{\prime}}i_{l}}A_{j_{l^{\prime}}i_{l}}-\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}[\boldsymbol{\gamma_{t,i}}]},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}.$
Similarly, if $j_{l^{\prime}}\neq j_{1}$ and $i_{l}\neq j_{1}$, then $[A_{j_{l^{\prime}}i_{l}},\boldsymbol{g}_{j}]=0$. Furthermore, remind that $\xi_{ji_{l}}=\xi_{i_{l}j},A_{ji_{l}}=A_{i_{l}j}$. Therefore,
@@
$\mathcal{P}_{1}^{j}$ $=\mathrm{i}\,\mathrm{t}\mathrm{r}\Big{(}\!\Big{[}\frac{1}{cN}\sum_{i_{l}\neq j_{1}}\big{(}\xi_{ji_{l}}-{w}_{ji}\big{)}A_{ji_{l}}+\frac{1}{cN}\sum_{i_{l}\neq j_{1}}{w}_{ji}A_{ji_{l}}\!-\!\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}[{\gamma_{t,i}}]},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}$
$=\mathrm{i}\,\underbrace{\mathrm{t}\mathrm{r}\Big{(}\!\Big{[}\frac{1}{cN}\sum_{i_{l}\neq j_{1}}\big{(}\xi_{ji_{l}}-{w}_{ji}\big{)}A_{ji_{l}},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}}_{(\star)^{j_{1}}}+\mathrm{i}\underbrace{\mathrm{t}\mathrm{r}\Big{(}\!\Big{[}\frac{1}{cN}\sum_{i_{l}\neq j_{1}}{w}_{ji}A_{ji_{l}}\!-\!\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}[{\gamma_{t,i}}]},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}}_{(\star\star)^{j_{1}}}$
@@
$\Big{|}\mathbb{E}[(\star)^{j_{1}}]\Big{|}$ $\leq\Big{|}\mathbb{E}\Big{[}\mathrm{t}\mathrm{r}\Big{(}\!\Big{[}\frac{1}{cN}\sum_{i_{l}\neq j_{1}}\big{(}\xi_{ji_{l}}-{w}_{ji}\big{)}A_{ji_{l}},\boldsymbol{g}_{t,j}\Big{]}\boldsymbol{\rho}_{t}^{cN}\Big{)}\Big{]}\Big{|}$
$\leq C\Big{|}\mathbb{E}\Big{[}\frac{1}{cN}\sum_{i_{l}\neq j_{1}}\Big{(}\xi_{ji_{l}}-{w}_{ji}\Big{)}\Big{]}\Big{|}$
@@
$\Big{|}\mathbb{E}[(\star)^{j_{1}}]\Big{|}$ $\leq C\Big{(}\Big{|}\frac{1}{cN}\sum_{i_{l}\neq j_{1}}\underbrace{\mathbb{E}\Big{[}\Big{(}\xi_{ji_{l}}-{w}^{G_{cN}}\big{(}\frac{jN-N+1}{cN},\frac{iN-N+l}{cN}\big{)}\Big{)}\Big{]}}_{0}\Big{|}$
$\quad\quad+\Big{|}\frac{1}{cN}\sum_{i_{l}\neq j_{1}}\mathbb{E}\Big{[}\Big{(}{w}^{G_{cN}}\big{(}\frac{jN-N+1}{cN},\frac{iN-N+l}{cN}\big{)}-{w}_{ji}\Big{)}\Big{]}\Big{|}\Big{)}.$
@@
$\frac{1}{c}\sum_{j=1}^{c}\big{|}\mathbb{E}[(\star)^{j_{1}}]\big{|}$ $=\frac{1}{cN}\sum_{j_{l^{\prime}}=1_{1}}^{c_{N}}\big{|}\mathbb{E}[(\star)^{j_{l^{\prime}}}]\big{|}$
$\leq C\Big{(}\Big{(}\frac{1}{cN}\Big{)}^{2}\sum_{j_{l^{\prime}}=1_{1}}^{c_{N}}\Bigl{|}\sum_{i_{l}=1_{1}}^{c_{N}}\mathbb{E}\Big{[}\Big{(}{w}^{G_{cN}}\big{(}\frac{jN-N+l^{\prime}}{cN},\frac{iN-N+l}{cN}\big{)}-{w}_{ji}\Big{]}\Bigr{|}\Big{)}$
@@
Estimate $(\star\star)^{j_{1}}$

### v3

@@ 5.2 Propagation of chaos
$\boldsymbol{\rho}_{0}^{cN}=\rho_{0}^{\otimes(cN)}\xrightarrow[]{\text{Propagation of chaos}}\boldsymbol{\rho}_{t}^{cN}\approx\bigotimes_{j=1}^{c}\bigotimes_{l=1}^{N}\gamma_{t,j_{l}},$
where for each class, we consider $N$ independent copies, each driven by its own Wiener process. Denote by $\gamma_{j_{l^{\prime}}}$ the $l^{\prime}$-th particle in class $j$, whose dynamics follow the SDE:
$\mathrm{d}\gamma_{t,j_{l^{\prime}}}$ $=-\mathrm{i}\Big[\tilde{H}+\frac{1}{c}\sum_{i=1}^{c}w_{ji}A^{\mathbb{E}_{\mathbb{P}}[\gamma_{t,i}]},\gamma_{t,j_{l^{\prime}}}\Big]\mathrm{d}t+\big(L\gamma_{t,j_{l^{\prime}}}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\gamma_{t,j_{l^{\prime}}}\}\big)\mathrm{d}t$
@@
To address the mean-field limit, we extend the method introduced by Pickl [66] and later adapted to the stochastic setting in [45, 47] to the case of a multi-class bosonic system.
The goal is to introduce a quantity that detects whether $\boldsymbol{\rho}^{cN}_{t}$ remains tensorized in a “good way” at each time. To this end, it suffices to show that the number of “bad” particles, namely particles deviating from the desired state $\boldsymbol{\gamma}_{i_{l}}$, is small.
To quantify the ratio of bad particles, we use a counting operator that runs through all possible combinations of bad particles in each class. We introduce the operator counting the number of bad particles in class $i$:
@@
$\mathbf{P}_{\hat{n}}$ $:=\prod_{i=1}^{c}\mathbf{P}_{i,n_{i}},\qquad\hat{n}=\begin{pmatrix}n_{1}\\ \vdots\\ n_{c}\end{pmatrix}$.
We can now define a functional that quantifies the number of bad particles in the system by weighting all possible combinations of bad particles.
For any sub-additive function $\varpi:\{0,\dots,cN\}\to\mathbb{R}_{+}$, we set
@@
which can equivalently be written as
@@
Moreover, if we count in an unbiased way, i.e., $\varpi(k)=\tfrac{k}{cN}$, then
$\mathbb{E}_{\mathbb{P}}\Big[\mathcal{E}_{N,\varpi}(t)\Big]\leq\frac{1}{c}\sum_{i=1}^{c}\mathbb{E}_{\mathbb{P}}\Big[\mathcal{D}^{i_{l}}_{N}(t)\Big],\quad\forall l\in[N],$
@@
$\stackrel{{\scriptstyle\text{Fubini}}}{{=}}\langle\boldsymbol{\Psi}_{t}^{cN},\sum_{i=1}^{c}\sum_{n_{1},\dots,n_{c}=0}^{N}\Big(\varpi(n_{i})\mathbf{P}_{i,n_{i}}\prod_{j\neq i}^{c}\mathbf{P}_{j,n_{j}}\Big)\boldsymbol{\Psi}_{t}^{cN}\rangle$
@@
$\mathbb{E}_{\mathbb{P}}\Big[\mathcal{E}_{N,\varpi}^{i}(t)\Big]$ $=\sum_{n_{i}=0}^{N}\mathbb{E}_{\mathbb{P}}\Big[\langle\boldsymbol{\Psi}_{t}^{cN},\frac{n_{i}}{cN}\mathbf{P}_{i,n_{i}}\boldsymbol{\Psi}_{t}^{cN}\rangle\Big],$
$=\frac{1}{c}\,\mathbb{E}_{\mathbb{P}}\Big[\langle\boldsymbol{\Psi}_{t}^{cN},\frac{1}{N}\sum_{n_{i}=0}^{N}{n_{i}}\mathbf{P}_{i,n_{i}}\boldsymbol{\Psi}_{t}^{cN}\rangle\Big],$
$=\frac{1}{c}\,\mathbb{E}_{\mathbb{P}}\Big[\langle\boldsymbol{\Psi}_{t}^{cN},\frac{1}{N}\sum_{n_{i}=0}^{N}\sum_{l=1}^{N}{\boldsymbol{g}_{t,i_{l}}}\mathbf{P}_{i,n_{i}}\boldsymbol{\Psi}_{t}^{cN}\rangle\Big],$
$\stackrel{{\scriptstyle\text{Invariance in law}}}{{=}}\frac{1}{c}\,\mathbb{E}_{\mathbb{P}}\Big[\langle\boldsymbol{\Psi}_{t}^{cN},\boldsymbol{g}_{t,i_{l}}\boldsymbol{\Psi}_{t}^{cN}\rangle\Big],\quad\forall l\in[N],$
$=\frac{1}{c}\mathbb{E}_{\mathbb{P}}\Big[\mathcal{D}_{N}^{i_{l}}(t)\Big],$
@@
The above lemma is crucial because it allows us to compute the number of bad particles class by class, and to work directly with the trace norm. Indeed, thanks to [44, Lemma 2.3], the quantity $\mathcal{D}_{N}^{i_{l}}(t)$ provides a control of the trace-norm distance between the reduced state of particle $i_{l}$ in the $cN$-particle system and the corresponding limiting state in the graphon system. More precisely, one has the bound
$\mathcal{D}^{i_{l}}_{N}(t)\leq\big\|\rho_{t,i_{l}}-\gamma_{t,i_{l}}\big\|_{1}\leq C\sqrt{\mathcal{D}^{i_{l}}_{N}(t)},$
for some universal constant $C>0$.
We now state some technical lemmas that are useful to establish our result. The following lemmas results from a slight modification of [47, Lemma A.1] and [44, Lemma 3.10].
Lemma 5.3.
Let $f:\mathbb{N}\to\mathbb{R}$ be a function and define the operator $\hat{f}_{i}=\sum_{k}f(k)\mathbf{P}_{i,k}$, where $\mathbf{P}_{i,k}$ is an operator counting the number of bad particles in class $i$.
Consider two products $\boldsymbol{Q}_{r_{i},r_{j}}^{(a)}$, $\boldsymbol{Q}_{r_{i},r_{j}}^{(b)}$ defined in the following way:
@@
$(\gamma_{2}$ $A_{12}\gamma_{2})f(x_{1},x_{2})=\int_{\mathfrak{X}}\gamma_{2}(x_{2},y_{2})(A_{12}\gamma_{2}f)(x_{1},y_{2})\mu(\mathrm{d}y_{2})$
@@
$\stackrel{{\scriptstyle\text{Fubini}}}{{=}}\int_{\mathfrak{X}^{2}}\Big(\int_{\mathfrak{X}^{2}}\gamma(x_{2},y_{2})a(x_{1},y_{2};x^{\prime}_{1},z_{2}){\gamma(z_{2},x^{\prime}_{2})}\mu(\mathrm{d}y_{2})\mu(\mathrm{d}z_{2})\Big)f(x_{1}^{\prime},x_{2}^{\prime})\mu(\mathrm{d}x_{1}^{\prime})\mu(\mathrm{d}x_{2}^{\prime})$
@@
The following theorem establishes the propagation of chaos.
Theorem 5.5 (Propagation of chaos estimate).
@@
$\mathbb{E}_{\mathbb{P}}[\mathcal{E}_{N,\hat{f}}(t)]$ $\leq e^{Ct}\Big(\mathcal{E}_{N,\hat{f}}(0)+\|{w}^{G_{cN}}-{w}\|_{\square}+\frac{1}{\sqrt{N}}\Big),$
@@
We proceed “à la Grönwall”. Recall that in unbiased counting, we have
$\mathbb{E}_{\mathbb{P}}[\mathcal{E}_{N,\hat{f}}(t)]\leq\frac{1}{c}\sum_{j=1}^{c}\mathbb{E}_{\mathbb{P}}[\mathcal{D}^{j_{l}}_{N}(t)],\quad\forall l\in[N].$
@@
where $\mathcal{P}_{1}^{j}$ collects the drift terms that are independent of $L_{j}:=L_{j_{1}}$ and $\mathcal{P}_{2}^{j}$ collects the drift terms that depend on $L_{j}$.
The estimates used to bound the quantity $|\mathcal{P}_{2}^{j}|$ are found in the lemmas of [45, Appendix A], and the extension for all density operators provided in [28]. We have
$\mathcal{P}_{2}^{j}\!$ $=\!\mathrm{t}\mathrm{r}\Big(\Big(\tfrac{1}{2}L_{j}^{\dagger}L_{j}\boldsymbol{\rho}_{t}^{cN}\!+\!\tfrac{1}{2}\boldsymbol{\rho}_{t}^{cN}L_{j}^{\dagger}L_{j}\!-\!L_{j}\boldsymbol{\rho}_{t}^{cN}L_{j}^{\dagger}\Big)\boldsymbol{\gamma_{t,j}}\!+\!\boldsymbol{\rho}_{t}^{cN}\Big(\tfrac{1}{2}L_{j}^{\dagger}L_{j}\boldsymbol{\gamma}_{t,j}\!+\!\tfrac{1}{2}\boldsymbol{\gamma}_{t,j}L^{\dagger}_{j}L_{j}\!-\!L_{j}\boldsymbol{\gamma}_{t,j}L_{j}^{\dagger}\Big)$
@@
$\mathcal{P}_{1}^{j}$ $=\mathrm{i}\,\mathrm{t}\mathrm{r}\Big(\boldsymbol{\rho}_{t}^{cN}[\tilde{\mathbf{H}}_{j},\boldsymbol{\gamma}_{t,j}]+\boldsymbol{\rho}_{t}^{cN}[\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}_{\mathbb{P}}[\boldsymbol{\gamma_{t,i}}]},\boldsymbol{\gamma}_{t,j}]\Big)$
@@
$\mathcal{P}_{1}^{j}$ $=\mathrm{i}\,\mathrm{t}\mathrm{r}\Big(\big[\tilde{\mathbf{H}}_{j},\boldsymbol{\gamma}_{t,j}\big]\boldsymbol{\rho}_{t}^{cN}+[\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}_{\mathbb{P}}[{\gamma_{t,i}}]},\boldsymbol{\gamma}_{t,j}]\boldsymbol{\rho}_{t}^{cN}\Big)$
@@
All terms with $\tilde{\mathbf{H}}_{i_{l}}$ vanish because $\tilde{\mathbf{H}}_{i_{l}},\boldsymbol{g}_{t,j}]=0$ when $i_{l}\neq j_{1}$  [^1] This follows since the operators $\mathbf{\tilde{H}}_{i_{l}},\boldsymbol{g}_{t,j}$ act on different variables. and the two remaining $\tilde{\mathbf{H}}_{j}$ terms cancel each other,
$\mathcal{P}_{1}^{j}$ $=\mathrm{i}\,\mathrm{t}\mathrm{r}\Big(\Big[\frac{1}{2cN}\sum_{i_{l}=1_{1}}^{c_{N}}\sum_{j_{l^{\prime}}=i_{l}}^{c_{N}}\xi_{j_{l^{\prime}}i_{l}}A_{j_{l^{\prime}}i_{l}},\boldsymbol{g}_{t,j}\Big]\boldsymbol{\rho}_{t}^{cN}\Big)-\mathrm{i}\,\mathrm{t}\mathrm{r}\Big(\Big[\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}_{\mathbb{P}}[\boldsymbol{\gamma_{t,i}}]},\boldsymbol{g}_{t,j}\Big]\boldsymbol{\rho}_{t}^{cN}\Big)$
$=\mathrm{i}\,\mathrm{t}\mathrm{r}\Big(\Big[\frac{1}{2cN}\sum_{i_{l}=1_{1}}^{c_{N}}\sum_{h_{k}=i_{l}}^{c_{N}}\xi_{j_{l^{\prime}}i_{l}}A_{j_{l^{\prime}}i_{l}}-\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}_{\mathbb{P}}[\boldsymbol{\gamma_{t,i}}]},\boldsymbol{g}_{t,j}\Big]\boldsymbol{\rho}_{t}^{cN}\Big).$
Similarly, if $j_{l^{\prime}}\neq j_{1}$ and $i_{l}\neq j_{1}$, then $A_{j_{l^{\prime}}i_{l}},\boldsymbol{g}_{j}]=0$. Furthermore, remind that $\xi_{ji_{l}}=\xi_{i_{l}j},A_{ji_{l}}=A_{i_{l}j}$. Therefore,
@@
$\mathcal{P}_{1}^{j}$ $=\mathrm{i}\,\mathrm{t}\mathrm{r}\Big(\!\Big[\frac{1}{cN}\sum_{i_{l}\neq j_{1}}\big(\xi_{ji_{l}}-{w}_{ji}\big)A_{ji_{l}}+\frac{1}{cN}\sum_{i_{l}\neq j_{1}}{w}_{ji}A_{ji_{l}}\!-\!\frac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}_{\mathbb{P}}[{\gamma_{t,i}}]},\boldsymbol{g}_{t,j}\Big]\boldsymbol{\rho}_{t}^{cN}\Big)$
$=\mathrm{i}\!$ $\underbrace{\mathrm{t}\mathrm{r}\Big(\!\Big[\tfrac{1}{cN}\sum_{i_{l}\neq j_{1}}\big(\xi_{ji_{l}}-{w}_{ji}\big)A_{ji_{l}},\boldsymbol{g}_{t,j}\Big]\boldsymbol{\rho}_{t}^{cN}\Big)}_{(\star)^{j_{1}}}\!+\mathrm{i}\!\underbrace{\mathrm{t}\mathrm{r}\Big(\!\Big[\tfrac{1}{cN}\sum_{i_{l}\neq j_{1}}{w}_{ji}A_{ji_{l}}\!-\!\tfrac{1}{c}\sum_{i=1}^{c}{w}_{ji}A_{j}^{\mathbb{E}_{\mathbb{P}}[{\gamma_{t,i}}]},\boldsymbol{g}_{t,j}\Big]\!\boldsymbol{\rho}_{t}^{cN}\Big)}_{(\star\star)^{j_{1}}}.$
@@
$\Big|\mathbb{E}_{\mathbb{P}}[(\star)^{j_{1}}]\Big|$ $\leq\Big|\mathbb{E}_{\mathbb{P}}\Big[\mathrm{t}\mathrm{r}\Big(\!\Big[\frac{1}{cN}\sum_{i_{l}\neq j_{1}}\big(\xi_{ji_{l}}-{w}_{ji}\big)A_{ji_{l}},\boldsymbol{g}_{t,j}\Big]\boldsymbol{\rho}_{t}^{cN}\Big)\Big]\Big|$
$\leq C\Big|\mathbb{E}_{\mathbb{P}}\Big[\frac{1}{cN}\sum_{i_{l}\neq j_{1}}\Big(\xi_{ji_{l}}-{w}_{ji}\Big)\Big]\Big|$
@@
$\Big|\mathbb{E}_{\mathbb{P}}[(\star)^{j_{1}}]\Big|$ $\leq C\Big(\Big|\frac{1}{cN}\sum_{i_{l}\neq j_{1}}\underbrace{\mathbb{E}_{\mathbb{P}}\Big[\Big(\xi_{ji_{l}}-{w}^{G_{cN}}\big(\frac{jN-N+1}{cN},\frac{iN-N+l}{cN}\big)\Big)\Big]}_{0}\Big|$
$\quad\quad+\Big|\frac{1}{cN}\sum_{i_{l}\neq j_{1}}\mathbb{E}_{\mathbb{P}}\Big[\Big({w}^{G_{cN}}\big(\frac{jN-N+1}{cN},\frac{iN-N+l}{cN}\big)-{w}_{ji}\Big)\Big]\Big|\Big).$
@@
$\frac{1}{c}\sum_{j=1}^{c}\big|\mathbb{E}_{\mathbb{P}}[(\star)^{j_{1}}]\big|$ $=\frac{1}{cN}\sum_{j_{l^{\prime}}=1_{1}}^{c_{N}}\Big|\mathbb{E}_{\mathbb{P}}[(\star)^{j_{l^{\prime}}}]\Big|$
$\leq C\Big(\big(\frac{1}{cN}\big)^{2}\sum_{j_{l^{\prime}}=1_{1}}^{c_{N}}\Big|\sum_{i_{l}=1_{1}}^{c_{N}}\mathbb{E}_{\mathbb{P}}\big[\Big(w^{G_{cN}}(\frac{jN-N+l^{\prime}}{cN},\frac{iN-N+l}{cN})-w_{ji}\big]\Big|$
@@
Estimate $(\star\star)^{j_{1}}$.